OpenStudy (anonymous):

Find the line parallel to y=-4x-1 going through (1,-3)

OpenStudy (anonymous):

You would use the same equation given keeping the same slope (y=-4x+b) and replace x and y with the coordinate (1,-3) to solve for b. After that you plug in b into the equation to get the answer.

OpenStudy (anonymous):

So basically your answer would be y=-4x+1

OpenStudy (anonymous):

Where do i plug B into?

OpenStudy (anonymous):

A linear equation is y=mx+b

OpenStudy (anonymous):

m equals slope and b is the y intercept

OpenStudy (anonymous):

Yes i know that

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

Where di you get the slope?

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

Never mind...

OpenStudy (anonymous):

First you would find the slope of the first equation, which is y = -4x-1 After you find the slope, the second line has to have the same slope as the first line. Use the equation y=mx + b, and solve for b. For x and y, plug in (1, -3)

OpenStudy (anonymous):

It says you are looking for a parallel line, there fore the slope must be congruent

OpenStudy (anonymous):

Both equations have to have the same slopes

OpenStudy (anonymous):

-3=-4(1)-1

OpenStudy (anonymous):

^ I plugged it in

OpenStudy (anonymous):

The slope of the first line is -4. y = mx + b -3 = -4(1) + b -3 = -4 + b b = 1. Plug in b = 1 to the equation y = mx + b, where m = -4. y = -4x + 1

OpenStudy (anonymous):

I get it! Thanks dude! <3

OpenStudy (anonymous):

No problem, glad you got the question.